A proton has a kinetic energy of 7.5 MeV. What is the speed of an alpha particle that has the same DeBroglie wavelength as the proton? The mass of the alpha particle is 6.65E-27 kg.

9517 m/s

9.53E6 m/s

1833 m/s

2.38E13 m/s

0.747 m/s
Answer #1

Answer :--

  • the de broglie wavelength of the proton is
  • \lambda _{p} = \frac{h}{p} = \frac{h}{\sqrt{2m_{p}K}}
  • \lambda _{p} = \frac{h}{\sqrt{2m_{p}K}}..........................(1)
  • now for alpha particle
  • \lambda _{\alpha} = \frac{h}{m_{ \alpha }v}................................(2)
  • but it is given that \lambda _{\alpha} = \lambda _{p}
  • so from above two equation
  • v = \frac{\sqrt{2m_{p}K_{p}}}{m_{\alpha}}
  • putting the values
  • K = 7.5 MeV = 1.2016 \times 10-12 J , mp = 1.673 \times 10-27 kg , malpha = 6.65 \times 10-27 kg
  • v = \frac{\sqrt{2\times 1.673 \times 10^{-27}\times 1.2016 \times 10^{-12}}}{6.65 \times10^{-27}}
  • v = 9.535 \times 10^{6} m/s
  • so option (B) is correct .